OpenStudy (anonymous):

True or false: cos(7x)-cos(5x)/sin(7x)-sin(5x)= tan(6x) is an identity what???

OpenStudy (anonymous):

Identity means "always true" and since the graph of these are literally mirrors of eachother it is not an identity.

OpenStudy (anonymous):

\[\cos C-\cos D=-2\sin \frac{ C+D }{ 2 }\sin \frac{ C-D }{ 2 }\] \[\sin C-\sin D=2 \cos \frac{ C+D }{ 2 }\sin \frac{ C-D }{ 2 }\]

OpenStudy (anonymous):

i am not sure what you used to get there

OpenStudy (anonymous):

\[\frac{ \cos C-\cos D }{ \sin C-\sin D }=-\tan \frac{ C+D }{ 2 }\] C=7 x,D=5x

OpenStudy (anonymous):

Which is 6, so its true. What formula is that?

OpenStudy (anonymous):

negative sign

OpenStudy (anonymous):

-tan 6x

OpenStudy (anonymous):

ahhhh so its false i missread that. That is exactly what i expected thank you.

OpenStudy (anonymous):

yw